# Planning Agents on an Ego-Trip: Leveraging Hybrid Ego-Graph Ensembles for Improved Tool Retrieval in Enterprise Task Planning

### v1

@@ Abstract
Effective tool retrieval is essential for AI agents to select from a vast array of tools when identifying and planning actions in the context of complex user queries. Despite its central role in planning, this aspect remains underexplored in the literature. Traditional approaches rely primarily on similarities between user queries and tool descriptions, which significantly limits retrieval accuracy, specifically when handling multi-step user requests. To address these limitations, we propose a Knowledge Graph (KG)-based tool retrieval framework that captures the semantic relationships between tools and their functional dependencies. Our retrieval algorithm leverages ensembles of 1-hop ego tool graphs to model direct and indirect connections between tools, enabling more comprehensive and contextual tool selection for multi-step tasks. We evaluate our approach on a synthetically generated internal dataset across six defined user classes, extending previous work on coherent dialogue synthesis and too retrieval benchmarks. Results demonstrate that our tool graph-based method achieves $91.85\%$ tool coverage on the micro-average CompleteRecall metric, compared to $89.26\%$ for re-ranked semantic-lexical hybrid retrieval, the strongest non-KG baseline in our experiments. These findings support our hypothesis that the structural information in the KG provides complementary signals to pure similarity matching, particularly for queries requiring sequential tool composition.
@@
Sahil Bansalthanks: *The authors contributed equally to this work., Sai Shruthi Sistla11footnotemark: 1, Aarti Arikatala, Sebastian Schreiber SAP Labs, Palo Alto {sahil.bansal01sai.shruthi.sistla aarti.arikatalasebastian.schreiber}@sap.com
@@ 1 Introduction
Agentic systems powered by Large Language (LLMs) or Reasoning Models (LRMs) excel in planning and scheduling sub-tasks for complex requests Kim et al. (2023); Erdogan et al. (2025); Rawat et al. (2025). While these systems effectively break down tasks into manageable logical sequences, evaluations have primarily focused on controlled settings with limited, well-defined tools that fit within a model’s context window, such as web search, a calculator, etc.
@@
Moreover, effective tool discovery is a prerequisite for meaningful task decomposition: agents must first identify what capabilities are available before they can decide how to break down and solve a complex problem. Despite its centrality to real-world agentic long-horizon planning, this aspect remains underexplored in the existing literature Huang et al. (2024); Wei et al. (2025).
We propose a structured semantic representation of enterprise tools using semi-structured data from tool descriptions and metadata. This approach creates a knowledge graph (KG) capturing relationships between tools, entities, and parameters, enabling better mapping of user queries to relevant tools. Our KG-enhanced mechanism uses neighborhood expansion to uncover implicit connections that traditional retrieval methods miss.
@@
    We introduce the Ensemble of Ego Graphs (EEG) algorithm, which uses an ensemble of 1-hop ego graphs extracted from our overall tool graph via a hybrid node matching and neighborhood expansion technique to improve tool retrieval performance.
@@
    We evaluate the retrieval efficacy of our EEG algorithm on complex user queries generated using the above pipeline, using a CompleteRecall metric Zhang et al. (2024) specifically adapted to our tool retrieval setup, demonstrating significant improvements over baseline approaches.
@@ 2 Related Work
Recent works have explored the use of tool graphs for tool retrieval, but exhibit notable limitations.
@@
For instance, ControlLLM Liu et al. (2024b) requires an adjacency matrix to construct the tool graph, assuming its structure is defined a priori. While offering a structured approach for tool selection and execution, this reliance on pre-specified connections limits its applicability in automation scenarios where tool relationships are unknown or evolving.
@@
ToolNet Liu et al. (2024a) employs graph-based iterative tool traversal similar to our approach. However, its graph construction depends on either extensive tool-use trajectories from code repositories and public datasets (unavailable for enterprise APIs) or LLM-generated trajectories, which frequently contain errors.
@@
COLT Qu et al. (2024) employs a complex multi-step, multi-bipartite graph training process for transductive graph embeddings, but lacks automatic scene inference and clear application paths for novel queries, limiting its generalization capabilities.
@@
Tool Graph Retriever Anonymous (2024) extracts tool dependencies from documentation to create a graph. Unlike their approach, our method doesn’t use a custom dependency identification model but instead leverages similarities between parameters and other entities, extracted via Open Information Extraction, to connect tools in a graph.
@@
This approach shares some similarities with ToolFlow Wang et al. (2025b), which also builds tool dependency graphs from documentation, but applies them to conversation generation rather than retrieval purposes. Building on the work presented in this paper, we further extended the methodology to synthetically generate complex, multi-step business queries in scenarios where output parameters are not available. This extension enables a more rigorous evaluation of the proposed graph-based retrieval mechanism.
@@
Graph RAG-Tool Fusion Lumer et al. (2025), developed concurrently with our work, similarly combines vector retrieval with knowledge graph traversal but differs in two key ways: First, while they use synthetic tool graphs with well-defined dependencies suited to depth-first search, our approach semi-automatically transforms real enterprise tools into graph representations. Second, we identify ego-graph entry nodes using multiple vector representations rather than limiting ourselves to only semantic embeddings, making our method more practical for enterprise environments.
@@
Another approach incorporates tool knowledge directly into model parameters through training or fine-tuning, including multi-label classifiers Moon et al. (2024) and LLMs Wang et al. (2025a). However, these parameter-based methods are inadequate for dynamic enterprise environments with large, frequently changing tool ecosystems. For a comprehensive review of such approaches and their limitations, we refer readers to recent surveys like Qu et al. (2025).
@@ 3.1 System Overview
Our proposed pipeline for tool retrieval is structured into two principal stages: an offline phase for building a structured KG from semi-structured tool documentation and metadata, and an online phase for retrieving relevant tools in response to user requests.
@@ 3.2 Knowledge Graph Construction
Table 1 presents a sample tool along with its metadata which will be referenced throughout the methodology section for better understanding.
Table 1: Sample tool with its metadata.
@@ 3.2.1 Ontology Definition
  • tool entity: Send Task Deadline Reminder to Team Members
  • With predicate-object pairs such as:
    • has_parameter $\rightarrow$ priority_filter
    • has_department $\rightarrow$ operations
By constraining both entity and relation types, we reduce graph noise, prevent combinatorial growth, and ensure that extracted triples directly support the system’s use-cases.
@@ 3.2.3 Triple Canonicalization
To ensure consistency and reduce redundancy in the KG, we perform normalization on both entities and predicates. This process involves two key steps:
@@
Algorithm 1 Knowledge Graph Construction
@@ 3.2.4 Graph Population
3.3 Ego Graph Retrieval
To retrieve the most relevant tools in response to a multi-step user query, as described in Algorithm 2, we employ a custom ego-graph retrieval algorithm consisting of three main stages:
3.3.1 Entry Point Identification in the Tool Graph
After identifying entry points, we execute a one-hop neighborhood expansion around each identified node, constructing an ensemble of ego graphs as previously described. This expansion process enriches our candidate set by incorporating all tool nodes directly connected to the entry points, thereby revealing contextually relevant tools that might otherwise remain undiscovered.
3.3.3 Re-Ranking Retrieved Tools
llama-3.2-nv-rerankqa-1b-v2222https://build.nvidia.com/nvidia/llama-3_2-nv-rerankqa-1b-v2/modelcard to re-rank the initially retrieved tools. The model takes as input the user query and each retrieved tool, and outputs a relevance score for each. We then retain the top-10 tools with the highest scores as the final output for a given user query.
Algorithm 2 Ego Graph Tool Retrieval
@@ 4 Dataset Generation
To evaluate our graph-based method, we require a tool retrieval benchmark suitable for enterprise use. Existing tool use benchmarks (Table 2) do not meet all these requirements, necessitating a custom dataset.
@@ 4.1 User Query Classification in Enterprise Task-Oriented Systems
  • Conditional Multi-Step Queries explicitly state a condition that must be met before executing the action.
@@ 4.2 Synthetic Multi-Step Query Generation
Here we introduce a structured Query Generation Pipeline covering all the aforementioned query types. This pipeline comprises three key components—Path Identification, Query Generation, and Query Validation—that collectively synthesize realistic, semantically grounded user queries spanning the various user query classes.
@@ 4.2.1 Path Identification
  • P-P Graph Construction: Inspired by ToolFlow Wang et al. (2025b), this graph captures semantic proximity between tools based on cosine similarity of input parameter embeddings. Each node corresponds to a tool, and edges are established when similarity exceeds a predefined threshold, suggesting potential sequential usage or shared functional behavior. This structure enables efficient exploration of tool compositions for multi-step planning.
  • Inferred R-P Graph Construction: We employ the o3-mini [^3] https://openai.com/index/openai-o3-mini/ reasoning model to infer plausible output parameters for individual tools. These outputs are matched to tools accepting them as inputs, forming directed output-to-input edges annotated with confidence scores. Each tool sequence is then passed through a validation step where we use o3-mini to reason where each generated sequence is valid. The prompts used can be found in the Appendix under section C. This graph reveals latent dependencies across tools, allowing for the construction of semantically valid but previously undocumented multi-tool flows.
@@ 4.2.2 Query Generation
This step operationalizes tool logic into realistic language patterns, enabling robust evaluation of retrieval systems under multi-step user query conditions.
@@ 4.3 Analysis of Generated Query Types
As a key outcome of our proposed pipeline, we generate a diverse set of synthetic queries spanning five complex query classes. We analyze the distribution of these queries and provide qualitative examples to showcase how the pipeline captures the structural and semantic characteristics of real-world enterprise queries. As seen in Figure 2, we categorized utterances into six distinct query classes based on intent category. The ‘single-intent’ category consists of real user queries sampled from production logs. The remaining query classes were synthetically generated to closely match this empirical distribution. Detailed examples for queries generated for each query type are present in Appendix B.1. This approach ensures our dataset reflects realistic usage patterns while enabling scalable coverage of more complex query types.
@@ 5.1 Experimental Setup
We constructed our graph database using an internal dataset comprising semi-structured information extracted from several hundred enterprise tools within a large software platform. This dataset includes detailed descriptions, parameter specifications, and associated metadata, serving as the foundational data for graph construction described in Section 3.2.
To evaluate our system, we employed a separate dataset of synthetic user queries generated through the pipeline outlined in Section 4.2.2. These queries were created by selecting targeted subsets of tools, formulating logical reasoning paths, and designing multi-step task-oriented queries. Each query was subsequently refined through a combination of automated validation and manual review to ensure high fidelity and practical alignment with real-world enterprise use cases.
@@ 5.2 Evaluation Metrics
where $Q$ is the evaluation query set and $\mathbf{1}$ is an indicator function that returns 1 if Recall@k for a given query $q$ is exactly $1.0$ - meaning all required tools for that query are present within the top-k retrieved items, with k being the selected rank cutoff. This metric is tailored to planning systems where task breakdown depends on retrieving a complete set of expected tools. A comparable notion of complete recall has been used in prior work in the context of table retrieval Zhang et al. (2024).
@@ 5.3 Retrieval Results and Analysis
To evaluate retrieval effectiveness, we compared four approaches: semantic retrieval, lexical retrieval (Okapi BM25444https://pypi.org/project/rank-bm25/), hybrid retrieval, as well as our proposed KG-based retrieval method. Table 3 summarizes the CompleteRecall metric across all four approaches broken down by query type as well as aggregated averages. Semantic retrieval using dense vector embeddings to capture the similarity between queries and tool descriptions achieved a micro-average CompleteRecall of $59.84\%$ at $k=3$, $73.96\%$ at $k=5$, and $85.69\%$ at $k=10$, but struggled with nuanced queries requiring deeper contextual understanding.
@@
Our hybrid baseline combines the top-10 results from both semantic and lexical retrieval approaches, re-ranking them using llama-3.2-nv-rerankqa-1b-v2. This strategy achieves micro-average CompleteRecall of $62.43\%$ at $k=3$, $78.93\%$ at $k=5$, and $89.26\%$ at $k=10$, outperforming the standalone methods.
@@
Our knowledge graph (KG)-based approach significantly outperformed all other methods, achieving micro-average CompleteRecall scores of $70.58\%$ at $k=3$, $79.13\%$ at $k=5$, and $91.85\%$ at $k=10$, as shown in Figure 3. Figure 5 highlights the most significant gains in complete recall within the conditional multi-step and implicit multi-step query categories. We attribute these improvements to the KG’s ability to explicitly model semantic relationships among tools, enabling context-aware retrieval and revealing connections through shared functionalities and data dependencies.

### v2

@@ Abstract
Effective tool pre-selection via retrieval is essential for AI agents to select from a vast array of tools when identifying and planning actions in the context of complex user queries. Despite its central role in planning, this aspect remains underexplored in the literature. Traditional approaches rely primarily on similarities between user queries and tool descriptions, which significantly limits retrieval accuracy, specifically when handling multi-step user requests. To address these limitations, we propose a Knowledge Graph (KG)-based tool retrieval framework that captures the semantic relationships between tools and their functional dependencies. Our retrieval algorithm leverages ensembles of 1-hop ego tool graphs to model direct and indirect connections between tools, enabling more comprehensive and contextual tool selection for multi-step tasks. We evaluate our approach on a synthetically generated internal dataset across six defined user classes, extending previous work on coherent dialogue synthesis and tool retrieval benchmarks. Results demonstrate that our tool graph-based method achieves $91.85\%$ tool coverage on the micro-average CompleteRecall metric, compared to $89.26\%$ for re-ranked semantic-lexical hybrid retrieval, the strongest non-KG baseline in our experiments. These findings support our hypothesis that the structural information modeled in the graph provides complementary signals to pure similarity matching, particularly for queries requiring sequential tool composition.
@@
Sahil Bansalthanks: The authors contributed equally to this work., Sai Shruthi Sistla11footnotemark: 1, Aarti Arikatala, Sebastian Schreiber SAP Labs, Palo Alto, CA, USA {sahil.bansal01, sai.shruthi.sistla aarti.arikatala, sebastian.schreiber}@sap.com
@@ 1 Introduction
Agentic systems powered by Large Language (LLMs) or Reasoning Models (LRMs) excel in planning and scheduling sub-tasks for complex requests kim2023llmcompiler; erdogan2025planandact; rawat2025preact. While these systems effectively break down tasks into manageable logical sequences, evaluations have primarily focused on controlled settings with limited, well-defined tools that fit within a model’s context window, such as web search, a calculator, etc.
@@
Moreover, effective tool discovery is an essential prerequisite for meaningful task decomposition: Agents must first identify what capabilities are available before they can decide how to break down and solve a complex problem. Despite its centrality to real-world agentic long-horizon planning, this aspect remains underexplored in the existing literature huang2024understanding; wei2025plangenllms.
We propose a structured semantic representation of enterprise tools using semi-structured data from tool descriptions and metadata. This approach produces a knowledge graph (KG) capturing relationships between tools, entities, and parameters, enabling better mapping of user queries to relevant tools. Our KG-enhanced retrieval mechanism uses neighborhood expansion to uncover implicit connections that traditional retrieval methods miss.
@@
    We introduce the Ensemble of Ego Graphs (EEG) algorithm, which uses an ensemble of 1-hop ego tool graphs extracted from our overall tool graph via a hybrid node matching and neighborhood expansion technique to improve tool retrieval performance.
@@
    We evaluate the retrieval efficacy of our EEG algorithm on complex user queries generated using the above pipeline, using a CompleteRecall metric zhang-etal-2025-murre specifically adapted to our tool retrieval setup, demonstrating significant improvements over baseline approaches.
@@ 2 Related Work
Recent works have explored the use of tool graphs for tool retrieval but exhibit notable limitations.
@@
For instance, ControlLLM liu2024controlllm requires an adjacency matrix to construct the tool graph, assuming its structure is defined a priori. While offering a structured approach for tool selection and execution, this reliance on pre-specified connections limits its applicability in automation scenarios where tool relationships are unknown or evolving.
@@
ToolNet liu2024toolnet employs graph-based iterative tool traversal similar to our approach. However, its graph construction depends on either extensive tool-use trajectories from code repositories and public datasets (unavailable for enterprise APIs) or LLM-generated trajectories, which frequently contain errors.
@@
COLT qu2024colt employs a complex multi-step, multi-bipartite graph training process for transductive graph embeddings, but lacks automatic scene inference and clear application paths for novel queries, limiting its generalization capabilities.
@@
Tool Graph Retriever anonymous2024tool extracts tool dependencies from documentation to create a graph. Unlike their approach, our method doesn’t use a custom dependency identification model but instead leverages similarities between parameters and other entities, extracted via Open Information Extraction, to connect tools in a graph.
@@
Our approach also shares some similarities with ToolFlow wang-etal-2025-toolflow, which builds tool dependency graphs from documentation as well, but applies them to conversation generation rather than retrieval purposes. Building on the work presented in this paper, we further extended the methodology to synthetically generate complex, multi-step business queries in scenarios where output parameters are not available. This extension enables a more rigorous evaluation of the proposed graph-based retrieval mechanism.
@@
Graph RAG-Tool Fusion lumer2025graphragtoolfusion, developed concurrently with our work, similarly combines vector retrieval with knowledge graph traversal but differs in two key ways: First, unlike their method that relies on synthetic tool graphs with well-defined dependencies suited to depth-first search, our approach semi-automatically converts real enterprise tools into graph representations, requiring only minimal manual input to tune prompts for accurate LLM interpretation of tool metadata and to define domain-specific ontologies and entity types that ensure semantic consistency during graph construction. Second, we identify ego-graph entry nodes using multiple vector representations rather than limiting ourselves to only semantic embeddings, making our method more practical for enterprise environments.
@@
Another related approach incorporates tool knowledge directly into model parameters through training or fine-tuning, including multi-label classifiers moon2024efficient and LLMs DBLP:conf/iclr/WangHJWB025. However, these parameter-based methods are inadequate for dynamic enterprise environments with large, frequently changing tool ecosystems. For a comprehensive review of such approaches and their limitations, we refer readers to recent surveys like Qu_2025.
@@ 3.1 System Overview
Our proposed pipeline for tool retrieval is structured into two principal stages: an offline phase for building a structured Knowledge Graph (KG) from semi-structured tool documentation and metadata, and an online phase for retrieving relevant tools in response to user requests.
@@ 3.2 Knowledge Graph Construction
Table 1 introduces an example tool and its metadata, which we use as a running example throughout the methodology section.
@@ 3.2.1 Ontology Definition
  • Entity:
    • type: tool
    • name: Send Task Deadline Reminder to Team Members
  • Predicate-object pairs, such as:
    • has_parameter: priority_filter
    • has_department: operations
By constraining both entity and relation types, we reduce graph noise, prevent combinatorial growth, and ensure that extracted triples directly support the system’s use-cases. The prompt which provides this guidance to the triple extraction LLM is shared in Figure 6.
@@ 3.2.2 Triple Extraction
While some variations in entity or relation phrasing remain, LLM-based extraction mitigates many of the weaknesses of traditional approaches (e.g., rigid pattern-matching, limited semantic generalization) as shown in recent surveys of relation extraction using large language models diaz2025survey; xu2024large. Residual inconsistencies are addressed through a triple canonicalization step, which we describe next.
@@ 3.2.3 Triple Canonicalization
To enhance consistency and reduce redundancy in the KG, we perform normalization on both entities and predicates. This process involves two key steps:
@@
Algorithm 1 Knowledge Graph Construction33footnotemark: 3
@@ 3.2.4 Graph Population
Algorithm 2 Ego Graph Tool Retrieval55footnotemark: 5
@@
3.3 Ego Graph Retrieval
To retrieve the most relevant tools in response to a single or multi-step user query, we employ a custom ego-graph retrieval algorithm, described in Algorithm 5, which consists of three main stages:
3.3.1 Entry Point Identification in the Tool Graph
After identifying entry points, we execute a one-hop neighborhood expansion around each identified node, constructing an ensemble of ego tool graphs as previously described. This expansion process enriches our candidate set by incorporating all tool nodes directly connected to the entry points, thereby revealing contextually relevant tools that might otherwise remain undiscovered.
3.3.3 Re-Ranking Retrieved Tools
The llama-3.2-nv-rerankqa-1b-v2666https://build.nvidia.com/nvidia/llama-3_2-nv-rerankqa-1b-v2/modelcard re-ranking model is used to re-rank the initially retrieved set of tools. The model takes as input the user query and each retrieved tool and outputs a relevance score for each. We retain the top-10 tools with the highest scores as the final output for a given user query.
An end-to-end example, along with a sample graph snippet, is provided in Appendix under section A.
@@ 4 Dataset Generation
To evaluate our graph-based method, we require a tool retrieval benchmark suitable for enterprise use. Existing tool use benchmarks do not meet all these requirements, necessitating a custom dataset, cp. Table 2.
@@ 4.1 User Query Classification in Enterprise Task-Oriented Systems
  • Conditional Multi-Step Queries explicitly state a condition that must be met before executing some of the actions involved.
@@ 4.2 Synthetic Multi-Step Query Generation
We introduce a structured Query Generation Pipeline covering all the aforementioned query types. This pipeline comprises three key components—Path Identification, Query Generation, and Query Validation—that collectively synthesize realistic and semantically grounded user queries spanning the various user query classes.
@@ 4.2.1 Path Identification
  • P-P Graph Construction: Inspired by ToolFlow wang-etal-2025-toolflow, this graph captures semantic proximity between tools based on cosine similarity of input parameter embeddings. Each node corresponds to a tool, and edges are established when similarity exceeds a predefined threshold, suggesting potential sequential usage or shared functional behavior. This structure enables efficient exploration of tool compositions for multi-step planning.
  • Inferred R-P Graph Construction: We employ the o3-mini [^7] https://openai.com/index/openai-o3-mini/ reasoning model to infer plausible output parameters for individual tools. These outputs are matched to tools accepting them as inputs, forming directed output-to-input edges annotated with confidence scores. Each tool sequence is passed through a validation step where we use o3-mini to reason where each generated sequence is valid. This graph reveals latent dependencies across tools, allowing for the construction of semantically valid but previously undocumented multi-tool flows.
@@ 4.2.2 Query Generation
This step transforms tool logic into realistic language patterns, enabling robust evaluation of retrieval systems under multi-step user query conditions.
@@ 4.2.3 Query Validation
Detailed example prompts for R-P graph construction, query creation and validation are shared in Figures 79 in Appendix C.
Query Type Example
@@ 4.3 Analysis of Generated Query Types
As a key outcome of our proposed pipeline, we generate a diverse set of synthetic queries spanning across five complex query classes. We analyze the distribution of these queries and provide qualitative examples to showcase how the pipeline captures the structural and semantic characteristics of real-world enterprise queries. As seen in Figure 2, we categorized utterances into six distinct query classes based on intent category. The single-intent category consists of real user queries sampled from production logs. The examples for the remaining query classes were synthetically generated using the method explained above to closely match this empirical distribution. Detailed examples for queries generated for each query type are present in Table 3. This approach ensures our dataset reflects realistic usage patterns while enabling scalable coverage of more complex query types.
@@ 5.1 Experimental Setup
We constructed our graph database using an internal dataset comprising semi-structured information extracted from several hundred enterprise tools within a large software platform. This dataset includes detailed descriptions, parameter specifications, and associated metadata, serving as the foundational data for the graph construction described in Section 3.2.
To evaluate our system, we employed a separate dataset of synthetic user queries generated through the pipeline outlined in Section 4.2. These queries were created by selecting targeted subsets of tools, formulating logical reasoning paths, and designing multi-step task-oriented queries. Each query was subsequently refined through a combination of automated validation and manual review to ensure high fidelity and practical alignment with real-world enterprise use cases.
@@ 5.2 Evaluation Metrics
where $Q$ is the evaluation query set and $\mathbf{1}$ is an indicator function that returns 1 if Recall@k for a given query $q$ is exactly 1.01.0—meaning all required tools for that query are present within the top-k retrieved items, with k being the selected rank cutoff. This metric is tailored to planning systems where task breakdown depends on retrieving a complete set of expected tools. A comparable notion of complete recall has been used in prior work in the context of table retrieval zhang-etal-2025-murre.
@@ 5.3 Retrieval Results and Analysis
To evaluate retrieval effectiveness, we compared four approaches: semantic retrieval, lexical retrieval (Okapi BM25888https://pypi.org/project/rank-bm25/), hybrid retrieval, as well as our proposed KG-based retrieval method. Table 4 summarizes the CompleteRecall metric across all four approaches broken down by query type as well as aggregated averages. Comparable graph-based baselines are not available because no existing work provides enterprise-grade tools or real-world enterprise setups of this kind. As a result, direct empirical comparison is not feasible.
Semantic retrieval using dense vector embeddings to capture the similarity between queries and tool descriptions achieved a micro-average CompleteRecall of $59.84\%$ at $k=3$, $73.96\%$ at $k=5$, and $85.69\%$ at $k=10$, but struggled with nuanced queries requiring deeper contextual understanding.
@@
Our hybrid baseline combines the top-10 results from both semantic and lexical retrieval approaches, and re-ranks them using llama-3.2-nv-rerankqa-1b-v2. This strategy achieves micro-average CompleteRecall of $62.43\%$ at $k=3$, $78.93\%$ at $k=5$, and $89.26\%$ at $k=10$, outperforming the standalone methods.
@@
Our knowledge graph-based approach significantly outperformed all other methods, achieving micro-average CompleteRecall scores of $70.58\%$ at $k=3$, $79.13\%$ at $k=5$, and $91.85\%$ at $k=10$, as shown in Figure 3. Figure 4 highlights the most significant gains in complete recall within the conditional multi-step and implicit multi-step query categories. Appendix B provides further insights into our observations. We attribute these improvements to the KG’s ability to explicitly model semantic relationships among tools, enabling context-aware retrieval and revealing connections through shared functionalities and data dependencies.
Figure 4: CompleteRecall@10 for each query type